OpenStudy (anonymous):

Solve: (2-3i)(5i)/ 2+3

OpenStudy (lalaly):

complex?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I don't think that's it. You have to use the conjugate of the bottom

OpenStudy (anonymous):

oh.

OpenStudy (lalaly):

\[\frac{(2 \times 5i)-(3i \times 5i)}{5}= \frac{10i - 15i^2}{5}\]\[i=\sqrt{-1} ... i^2=-1\]\[\frac{10i-15(-1)}{5}= \frac{10i+15}{5} = 2i+3\]

OpenStudy (anonymous):

Distibute the top, and then deal with the bottom. There isn't any subtraction sign in the top part

OpenStudy (lalaly):

the bottom is 2+3 which is 5 and i already distributed whats in the numerator

OpenStudy (anonymous):

2 plus 3 "I"

OpenStudy (lalaly):

oh u forgot the i

OpenStudy (anonymous):

Sorry :O

OpenStudy (anonymous):

Why 2-9?

OpenStudy (lalaly):

3i*3i = -9 in the denominator

OpenStudy (anonymous):

on the bottom I got 4-9i and then that =13

OpenStudy (anonymous):

and why is 20i not -20i

OpenStudy (lalaly):

yeah that was my mistake im sorry ... difference between two squares (2-3i)(2+3i)=4-(-9)

OpenStudy (lalaly):

wait let me do it again

OpenStudy (anonymous):

yes, it is a positive 20, my bad :D

OpenStudy (anonymous):

wait no, take back!

OpenStudy (lalaly):

\[\frac{10i+15}{2+3i} \times \frac{2-3i}{2-3i} = \frac{20i-30(-1)+30-45i}{4-(-9)}\]\[=\frac{-25i+60}{13}\]

OpenStudy (anonymous):

When you multiply the -10 and the 2i should it be -20i

OpenStudy (lalaly):

what -10 and 2i?

OpenStudy (lalaly):

thers 10i and 2

OpenStudy (anonymous):

but the 10 is negative?

OpenStudy (lalaly):

no when u distribute at the begining 5i*2 = 10i

OpenStudy (anonymous):

I am going to attach my work

OpenStudy (lalaly):

kk

OpenStudy (anonymous):

can in progress. sorry

OpenStudy (anonymous):

*scan

OpenStudy (lalaly):

well the solution for \[\frac{(2-3i)(5i)}{2+3i}\]is what i did , last thing

OpenStudy (anonymous):

Thanks